OpenStudy (anonymous):

hi can anyone help me with this problem? (a^2+5a+4/a^3)/(a^2+3a+2/a^2-2a) please show your work so i know what youre doing,,

OpenStudy (anonymous):

if it helps written out it would be \[\frac{ a ^{2} + 5a+4 }{ a ^{3} } \div \frac{ a ^{2}+3a+2 }{ a ^{2}-2a }\]

OpenStudy (freckles):

do you know how to factor any of the expressions here?

OpenStudy (perl):

multiply by the reciprocal and factor

OpenStudy (anonymous):

how would i factor them

OpenStudy (perl):

$$\frac{ a ^{2} + 5a+4 }{ a ^{3} } \div \frac{ a ^{2}+3a+2 }{ a ^{2}-2a } \\~\\ =\frac{ (a+1)(a+4)}{ a ^{3} } \div \frac{ (a+1)(a+2) }{a(a-2) } \\~\\ =\frac{ (a+1)(a+4)}{ a ^{3} } \cdot \frac{ a(a-2) }{ (a+1)(a+2) } $$

OpenStudy (anonymous):

thanks, would I subtract out the a+1 ? or no

OpenStudy (perl):

yes, cancel out

OpenStudy (anonymous):

could I also cancel out the a on the numerator and denominator of the second fraction?

OpenStudy (perl):

yes :)

OpenStudy (anonymous):

is this right \[\frac{ a+4 }{ a ^{3} } \div \frac{ a-2 }{ 2 }\]

OpenStudy (perl):

shouldn't be a division sign

OpenStudy (anonymous):

oh right sorry i knew that

OpenStudy (anonymous):

but is that right

OpenStudy (perl):

$$ \frac{ a ^{2} + 5a+4 }{ a ^{3} } \div \frac{ a ^{2}+3a+2 }{ a ^{2}-2a } \\~\\ =\frac{ (a+1)(a+4)}{ a ^{3} } \div \frac{ (a+1)(a+2) }{a(a-2) } \\~\\ =\frac{ (a+1)(a+4)}{ a ^{3} } \cdot \frac{ a(a-2) }{ (a+1)(a+2) } \\~\\ =\frac{\cancel{ (a+1)}(a+4)}{ \cancel{a ^3} } \cdot \frac{ \cancel{a}(a-2) }{\cancel{(a+1)}(a+2) } \\~\\ =\frac{ (a+4)}{ a ^2 } \cdot \frac{ (a-2) }{ (a+2) } \\~\\ =\frac{ (a+4)(a-2)}{ a ^2(a+2) } $$

OpenStudy (anonymous):

and would the answer be \[\frac{ (a+4)(a-2) }{ a ^{2}(a+2) }\]

OpenStudy (anonymous):

I went with that and that was the right answer :)